OpenStudy (trojanpoem):

Circuits:

OpenStudy (trojanpoem):

I got the answer with KCL as -2.59 mA while the book answer is -3.41 mA

OpenStudy (trojanpoem):

Post my solution ?

OpenStudy (trojanpoem):

Req = 20/39 kohm Ieq = 4 mA V = 80/39 v I1 = 3 + IA 80/ (39 * 5) = 3 + IA Ia = -2.59 mA if I put -80/(39 * 5) I get the book answer, but I don't know what's wrong with the direction of I.

OpenStudy (trojanpoem):



OpenStudy (retireed):

I still looking at this but I am very rusty.

OpenStudy (retireed):